What is the length of a diagonal in a 78 foot long 27 foot wide rectangle?

Answer by checkley79(3341)

A^2+B^2=C^2
78^2+27^2=C^2
6,084+729=C^2
6,813=C^2
C=SQRT6,813
C=82.54 ANS. FOR THE DIAGONAL.